OpenStudy (anonymous):

Rationalize the denominator of square root of negative 36 over open parentheses 2 minus 3 i close parentheses plus open parentheses 3 plus 2 i

OpenStudy (anonymous):

start by simplifying the denominator expression: (2 - 3i) + (3 + 2i) = 2 + 3 - 3i + 2i = 5 - i Does that make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

square root of -36 gives 6i on the top. But to rationalize the bottom, you have to multiply by the complex conjugate... so since we found (5 - i) on the bottom, you multiply top and bottom by (5 + i). Then, there's a bunch of simplifying at the end.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

simplifying that last expression... = -6/26 + 30i/26 = -3/13 + (15/13)i Looks weird, but that's what I got. Without seeing your work, I'm not sure where we did it differently.

OpenStudy (anonymous):

thats fine with me

OpenStudy (anonymous):

At the end, though, you want the answer in the form of a number plus or minus some quantity with "i"... that is the form A + Bi, where A is the real number part and Bi is the imaginary part.

OpenStudy (anonymous):

soooo... i hope I helped

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np